How many kw-hr are consumed by twenty-five 60 watts light bulbs burning 5 hrs. in a 12v 

circuit?


1
Expert's answer
2021-09-17T15:33:07-0400

The total power dissipated on the bulbs

"P=NP_0=25*60\:\rm W=1500\: W"

The amount of electricity

"W=Pt\\
=1500\:\rm W*5\: hr=7500\: W\cdot hr=7.5\: kW\cdot hr"